# Classical simulability of constant-depth linear-optical circuits with noise

### v1

@@ V Partial distinguishability noise
where $\hat{\rho}_{I}$ is the state whose $I$ elements are indistinguishable and others are distinguishable and $p_{k}\equiv x^{k}(1-x)^{N-k}\binom{N}{k}^{-1}$. Then, the quantum state of the $N$ partially distinguishable single photons is equivalent to the mixture of an $N$ particle state, which is obtained by randomly selecting $k$ particles following a binomial distribution with success probability $x$ and setting them indistinguishable bosons and other fully distinguishable particles. Therefore, the remaining $N-k$ photons do not interfere with other particles. From the perspective of bipartite graphs and percolation, it corresponds to decoupling the corresponding vertices from the original bipartite graph as illustrated in Fig. 2. A difference from the photon loss effect is that whereas we remove the corresponding vertices for the case of photon loss, for the partial distinguishability effect, we remove the vertices and then construct bipartite graphs with each removed vertex. The percolation lemma still applies here because the resultant bipartite graphs still have the maximum size $O(\log(N/\epsilon))$ with high probability. Thus, the classical algorithm needs to be modified to simulate the systems corresponding to the new bipartite graphs with a single vertex on the left-hand side, and so we again have a similar theorem:
@@
For a given partial distinguishability $1-x$ and a linear-optical circuit of an arbitrary architecture of maximum degree $\Delta$ with $N$ single photon input, if $x<1/\Delta^{2}$, there exists a classical algorithm that can approximately simulate the corresponding lossy linear-optical circuits in $\text{poly}(N,1/\epsilon)$ within total variation distance $\epsilon$.
Figure 3: Largest size of the component obtained by removing vertices in $A$ of initial size $N$ with probability $1-\eta$. The upper panel’s initial bipartite graphs are obtained by randomly generating $\Delta$ edges from each vertex in $A$ to $M=8N$ number of vertices in $B$. The lower panel’s initial bipartite graphs are obtained by generating edges from each vertex in $A$ to $\Delta$ nearest vertices in $B$.
VI Numerical result
To clearly see the effect of photon loss or partial distinguishability on linear-optical systems, we numerically investigate the maximum size of components $\max_{i}|G_{i}|$ after randomly removing vertices from $A$. We consider two different cases: linear-optical circuits (i) with non-local beam splitters and (ii) with 1D-local beam splitters. For simplicity of the numerical simulation, in the non-local case, for a given input photon number $N$ and a number of modes $M$, we randomly generate $\Delta$, and, in the 1D case, we generate $\Delta$ edges from each input mode to the $\Delta$ closest output modes.
For a fixed $\Delta$ and different loss rates, we increase the number of input modes $N$ and analyze the largest component size, i.e., $\max_{i}|G_{i}|$, which is showcased in Fig. 3. First of all, it is clear that depending on the loss rates, the largest component size increases in distinct ways; when the loss rate is sufficiently low, it increases linearly, while the loss rate is high enough, it starts to increase logarithmically as expected by Lemma 1. We also emphasize that the transition point of the loss rate depends on the architecture. For instance, when $\Delta=9$, the transition occurs between ${\eta=0.4}$ and ${\eta=0.7}$ for 1D systems, whereas it occurs between ${\eta=0.02}$ and ${\eta=0.14}$, which implies that the non-local structure is much more robust than the 1D structure. Therefore, while Lemma 1 presents the worst-case threshold, the actual threshold may depend on the details of the architecture, such as the geometry of the architecture (e.g., non-local beam splitters or geometrically local beam splitters.) and how many output modes are relevant.
VII Discussion on more general cases
@@
Therefore, we can follow the same procedure as single-photon cases. A difference is that to see the percolation effect, now we sample from a binomial distribution with failure probability $(1-\eta)^{n}$, which corresponds to vacuum input and thus the associated vertex is removed from the bipartite graph. Consequently, the percolation threshold becomes
$[1-(1-\eta)^{n}]\Delta^{2}<1.$ (14)
For this case, since each input mode has at most $n$ photons, the associated Hilbert space’s dimension for $y^{*}$ is at most
@@
Similarly, for other more general input states than Fock states, if the largest photon number of each input state for each mode is bounded by a constant, then the Hilbert space’s dimension is at most polynomial in $N/\epsilon$; more generally, as long as the largest total number of photons is bounded by a linear function of $y^{*}$, the Hilbert space’s dimension is still upper bounded by $\text{poly}(N/\epsilon)$. For larger depth $d=\omega(1)$, we again need to use the matrix produce state method (see Appendix B for more details).
Hence, one may see that a sufficient condition for the percolation result to apply is that the lossy input state is written as
$\hat{\rho}=(1-p)|0\rangle\langle 0|+p\hat{\sigma},$ (16)
@@
VIII Discussion
In this work, we showed that a threshold of loss or noise rate exists above which classical computers can efficiently simulate constant-depth linear-optical circuits with certain input states. Our result implies that shallow-depth circuits may also be vulnerable to loss and noise because the entanglement in the system constructed by shallow-depth circuits may be more easily annihilated by noise.
An interesting future work is to find a general condition for the input states under which the percolation result gives the easiness result. Whereas Fock states’ threshold is easily found, it is not immediately clear to analytically find the threshold of more general quantum states, such as Gaussian states. Also, while our results hold for arbitrary architecture as long as the depth is constant, the depth limit might be pushed further depending on the details of the architecture, such as the geometry of the circuits and the input state configuration [46, 47]. Conversely, investigating the possibility of the hardness of constant-depth boson sampling with a loss rate below the threshold is another interesting future work enabling us to demonstrate quantum advantage even under practical loss and noise effects. Finally, we can easily see that the percolation lemma immediately applies for a continuous-variable erasure channel considered in Refs. [48, 49, 50]. Thus, we may be able to apply a similar technique to other noise models, such as more general Gaussian noise [51].
Acknowledgements.
We thank Byeongseon Go and Senrui Chen for interesting and fruitful discussions. This research was supported by Quantum Technology R&D Leading Program (Quantum Computing) (RS-2024-00431768) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT (MSIT))
Appendix A Proof of Lemma 1
In this Appendix, we provide the proof of Lemma  1 in the main text. The proof is based on Refs. [37, 38, 36] and is adapted to bipartite-graph cases.
@@ Appendix A Proof of Lemma 1
We present an algorithm that constructs a random graph $G^{\prime}=(A^{\prime},B^{\prime},E^{\prime})$ as described in the lemma. Denote the set of the vertices on the left-hand side as $A$ and on the left-hand side as $B$. We initialize this to be the empty graph. Then, we construct $S$ by querying the vertices on $A$. A query to vertex $v\in A$ succeeds with probability $\eta$, in which case the vertex is added to $S$. When $S=\emptyset$, the algorithm initializes $S$ by querying all unqueried vertices in $G$ until the first successful query. When $S\neq\emptyset$, the algorithm queries all unqueried vertices in $N_{G}(S)$, where $N_{G}(S)$ is a subset of $A$ which is connected by $S$ through $B$ by one step. The size of $N_{G}(S)$ is upper-bounded by $|S|\Delta^{2}$. Whenever the algorithm runs out of unqueried vertices, it adds $S$, the vertices in $B$ connected to $S$, and corresponding edges to $G^{\prime}$ and resets $S$ and continues. The algorithm finishes when there are no more unqueried vertices in $A$. Note that when the algorithm adds $S$ and its corresponding vertices in $B$ and edges to $G^{\prime}$, the added graph is always disconnected to the one in every step, which results in the set of disjoint components $\{G_{i}\}_{i=1}^{m}$, where $m$ is the number of steps.
@@
Appendix B Matrix product state for more general states than single photons
In this Appendix, we show that any linear-optical circuits with $N$ input states that have a constant maximum photon number can be simulated by matrix product state with bond dimension at most $c^{N}$; thus, the computational cost is exponential in $N$ with a constant $c$.
@@ Appendix B Matrix product state for more general states than single photons
Appendix C Upper bound of total variation distance
In this Appendix, we derive the upper bound of the approximation error caused by skipping the case where the connected component’s size is larger than $y^{*}$. Recall that the output probability $q$ of such an algorithm is given by
$q(\bm{m})=p(\bm{m}|E).$ (29)
$p(\bm{m})=p(\bm{m}|E)p(E)+p(\bm{m}|E^{\perp})p(E^{\perp}),$ (30)
$\sum_{\bm{m}}|p(\bm{m})-q(\bm{m})|$ (31)
$=\sum_{\bm{m}}p(\bm{m}|E)|p(E)-1|+p(E^{\perp})$ (34)
$=2p(E^{\perp})\leq 2\epsilon.$ (35)

### v2

@@ II.4 Partial distinguishability noise
where $\hat{\rho}_{I}$ is the state whose $I$ elements are indistinguishable and others are distinguishable and $p_{k}\equiv x^{k}(1-x)^{N-k}$. Then, the quantum state of $N$ partially distinguishable single photons is equivalent to the mixture of an $N$-particle state obtained by randomly selecting $k$ particles following a binomial distribution with success probability $x$ and setting them indistinguishable bosons and other fully distinguishable particles. Therefore, the remaining $N-k$ photons do not interfere with others. From the graph perspective, it corresponds to decoupling the corresponding vertices from the original graph as illustrated in Fig. 2. A difference from photon loss is that we construct bipartite graphs with each removed vertex. Lemma 1 still applies here because the resultant bipartite graphs still have the maximum size $O(\log(N/\epsilon))$ with high probability. Thus,
@@
For partial distinguishability $1-x$ and a linear-optical circuit of maximum degree $\Delta$ with $N$ single-photon input, if $x<1/\Delta^{2}$, there exists a classical algorithm that can approximately simulate the circuit in $\text{poly}(N,1/\epsilon)$ within TVD $\epsilon$.
Figure 3: Largest size of the component obtained by removing vertices in $A$ of initial size $N$ with probability $1-\eta$. The dashed lines indicate the size of graphs with 56 vertices, corresponding to the largest permanent computed so far using a supercomputer [64], i.e., the boundary of our classical algorithm. (See the main text for the details and Supplemental Material [35] for different scales of each axis.)
II.5 Numerical result
To clearly see the percolation effect, we numerically investigate the maximum size of components $\max_{i}|G_{i}|$ after randomly removing vertices from $A$. We consider two cases: linear-optical circuits (i) with non-local beam splitters and (ii) with 1D-local beam splitters. For simplicity of the numerical simulation, in the non-local case, for a given input photon number $N$ and the number of modes $M=8N$, we randomly generate $\Delta$ edges from each input, and, in the 1D case, we generate $\Delta$ edges from each input mode to the $\Delta$ closest output modes.
For a fixed $\Delta$ and different loss rates, we increase the number of input modes $N$ and analyze the largest component size, $\max_{i}|G_{i}|$, which is showcased in Fig. 3. Clearly, the largest component size increases in distinct ways depending on the loss rates; when the loss rate is sufficiently low, it increases linearly, while the loss rate is high enough, it starts to increase logarithmically as expected by Lemma 1 (See SM Sec. S2 [35].). We also emphasize that the transition point depends on the architecture. For instance, when $\Delta=9$, the transition occurs between ${\eta=0.4}$ and ${\eta=0.7}$ for 1D systems, whereas it occurs between ${\eta=0.02}$ and ${\eta=0.14}$, implying that the non-local structure is much more robust than the 1D structure. Therefore, while Lemma 1 presents the worst-case threshold, the actual threshold may depend on the details of the architecture, such as the geometry of the architecture (e.g., non-local beam splitters or geometrically local beam splitters.) and how many output modes are relevant.
Finally, we note that the state-of-the-art (Gaussian) boson sampling experiments suffer from a loss rate around $0.5-0.7$, i.e., $\eta\approx 0.3-0.5$ [9, 10, 11, 12] and that the largest single-photon boson sampling so far, to the best of our knowledge, is up to 20 photons [65]. Figure 3 shows that while a very shallow-depth nonlocal circuit with $\Delta=4$ is easy to simulate using our algorithm around this regime even when the initial number of photons is as large as $10^{4}$, as the circuit depth increases (e.g., $\Delta=9$), our algorithm starts to require a computational cost beyond the currently available supercomputers. A similar behavior can be observed for 1D linear-optical circuits. Overall, unless the loss rate is very low, around an initial hundred photons are required for shallow-depth boson sampling to be beyond our classical algorithm. It is worth emphasizing that this boundary does not necessarily imply the boundary of classical algorithms because there may exist better classical algorithms than the presented one.
II.6 Discussion on more general cases
@@
Therefore, we can follow the same procedure as single-photon cases. A difference is that we now sample from a binomial distribution with failure probability $(1-\eta)^{n}$, which corresponds to vacuum input, and thus, the associated vertex is removed from the bipartite graph. Consequently, the percolation threshold becomes ${[1-(1-\eta)^{n}]\Delta^{2}<1}$. For this case, since each input mode has at most $n$ photons, the associated Hilbert space’s dimension for $y^{*}$ is at most
@@
Similarly, for more general input states than Fock states, if the largest photon number of each input state for each mode is constant, the Hilbert space’s dimension is at most polynomial in $N/\epsilon$; more generally, as long as the largest total photon number is bounded by a linear function of $y^{*}$, the Hilbert space’s dimension is still upper-bounded by $\text{poly}(N/\epsilon)$. For larger depth $d=\omega(1)$, we again need to use the MPS method (see Methods).
@@
III Discussion
We showed that a threshold of loss or noise rate exists above which classical computers can efficiently simulate constant-depth linear-optical circuits with certain input states. Thus, shallow-depth circuits may also be vulnerable to loss and noise because the entanglement in the system constructed by shallow-depth circuits may be more easily annihilated by noise.
An interesting future work is to find a general condition for input states under which the percolation result gives the easiness result. Whereas Fock states’ threshold is easily found, it is not immediately clear to analytically find the threshold of more general quantum states, such as Gaussian states [66]. Also, while our results hold for arbitrary architecture as long as the depth is constant, the depth limit might be pushed further depending on the details of the architecture, such as the geometry of the circuits and the input state configuration [67, 68, 69]. Conversely, investigating the possibility of the hardness of constant-depth boson sampling with a loss rate below the threshold is another interesting future work enabling us to demonstrate quantum advantage even under practical loss and noise effects. Finally, we can easily see that the percolation lemma immediately applies for a continuous-variable erasure channel considered in Refs. [70, 71, 72]. Thus, we may be able to apply a similar technique to other noise models, such as more general Gaussian noise [66].
Methods
III.1 Proof of Lemma 1
We provide the proof of Lemma 1 in the main text. The proof is based on Refs. [55, 56, 54] and is adapted to bipartite-graph cases.
@@ III.1 Proof of Lemma 1
We present an algorithm that constructs a random graph $G^{\prime}=(A^{\prime},B^{\prime},E^{\prime})$ as described in the lemma. Denote the set of the vertices on the left-hand side as $A$ and on the right-hand side as $B$. We initialize this to be the empty graph. Then, we construct $S$ by querying the vertices on $A$. A query to vertex $v\in A$ succeeds with probability $\eta$, in which case the vertex is added to $S$. When $S=\emptyset$, the algorithm initializes $S$ by querying all unqueried vertices in $G$ until the first successful query. When $S\neq\emptyset$, the algorithm queries all unqueried vertices in $N_{G}(S)$, where $N_{G}(S)$ is a subset of $A$ which is connected by $S$ through $B$ by one step. The size of $N_{G}(S)$ is upper-bounded by $|S|\Delta^{2}$. Whenever the algorithm runs out of unqueried vertices, it adds $S$, the vertices in $B$ connected to $S$, and corresponding edges to $G^{\prime}$ and resets $S$ and continues. The algorithm finishes when there are no more unqueried vertices in $A$. Note that when the algorithm adds $S$ and its corresponding vertices in $B$ and edges to $G^{\prime}$, the added graph is always disconnected to the one in every step, which results in the set of disjoint components $\{G_{i}\}_{i=1}^{m}$, where $m$ is the number of steps.
@@
It is worth emphasizing that the lemma can be understood from a previous result [56, 54] by defining a graph based on the bipartite graph in such a way that the graph’s vertices are $A$ and two vertices have an edge if they are connected by a vertex in $B$ in the original bipartite graph. Physically, two input modes are connected if the photons from them can be outputted in the same output mode, i.e., they can interfere with each other. Using the relation between the underlying bipartite and the new graph $G$, we can easily see that the maximum degree of the graph $G$ is upper-bounded by $\Delta^{2}$.
III.2 Upper bound of total variation distance
We derive the upper bound of the approximation error caused by skipping the case where the connected component’s size is larger than $y^{*}$. Recall that the output probability $q$ of such an algorithm is given by
$q(\bm{m})=p(\bm{m}|E).$ (21)
$p(\bm{m})=p(\bm{m}|E)p(E)+p(\bm{m}|E^{\perp})p(E^{\perp}),$ (22)
$\sum_{\bm{m}}|p(\bm{m})-q(\bm{m})|$ (23)
$=\sum_{\bm{m}}p(\bm{m}|E)|p(E)-1|+p(E^{\perp})$ (26)
$=2p(E^{\perp})\leq 2\epsilon.$ (27)
III.3 Matrix product state for more general states than single photons
In this section, we show that any linear-optical circuits with $N$ input states that have a constant maximum photon number can be simulated by matrix product state with bond dimension at most $c^{N}$; thus, the computational cost is exponential in $N$ with a constant $c$.
@@ III.3 Matrix product state for more general states than single photons
Acknowledgements.
We thank Byeongseon Go and Senrui Chen for interesting and fruitful discussions. This research was supported by the National Research Foundation of Korea (NRF) Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)).
Author contributions
C.O. developed the theory, implemented the numerical experiments, and wrote the paper.
Competing interests
The author declares no competing interests.